Aqueous hydrochloric acid HCl will react with solid sodium hydroxide NaOH to produce aqueous sodium chloride NaCl and liquid water H2O. Suppose 3.28 g of hydrochloric acid is mixed with 2.2 g of sodium hydroxide. Calculate the maximum mass of Sodium Chloride that could be produced by the chemical reaction. Round your answer to 2 significant digits.


1
Expert's answer
2021-05-11T05:57:29-0400

given reaction is-



"HCl(aq)+NaOH(s)\to NaCl(aq)+H_2O(l)"

moles of HCl taken =


"=\frac{3.28}{36.5}=0.090\ mole"

"\frac{}{}" moles of NaOH taken


"=\frac{2.2}{40}=0.055\ mole"

So NaOH is the limiting reagent here

then

mass of Nacl formed = moles of Nacl multiplied by its molar mass

Moles of Nacl will be equal to moles of NaOH taken as they react in the ratio of 1:1.



"=0.055\times 58.44=3.2\ gm\ Nacl"